OpenStudy (anonymous):

last question.. promise :) A ball is tossed from a window that is 10 feet off the ground. The height, h (in feet), of the ball at any time, t (in seconds), is given by the equation h = - 16t^2 + 12x + 10. What is the maximum height (in feet) that the ball will reach? Type the number only in the box, do not put measurement units.

OpenStudy (anonymous):

guess what? it is once again \(-\frac{b}{2a}\) this time \[-\frac{12}{2\times (-16)}\]

OpenStudy (anonymous):

your fraction reduces to \(\frac{3}{8}\)

OpenStudy (anonymous):

do i solve after that?

OpenStudy (anonymous):

replace \(x\) by \(\frac{3}{8}\)

OpenStudy (mimi_x3):

don't you differentiate it for the maximum?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

our miss brooks is not taking calc, she is taking algebra

OpenStudy (anonymous):

i thought u squared the denominator?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[-\frac{b}{2a}\] there is no square

OpenStudy (anonymous):

so for max i just put 3/8

OpenStudy (anonymous):

replace \(t\) by \(\frac{3}{8}\) get \(\frac{49}{4}\) as your answer

OpenStudy (anonymous):

12.25?

OpenStudy (anonymous):

\(\frac{3}{8}\)is not the max, it is the number that you use to find the max

OpenStudy (anonymous):

yes, 12.25

OpenStudy (anonymous):

yayyy thanks sooo much :)

OpenStudy (anonymous):

you done?

OpenStudy (anonymous):

yes i just needed the max